OpenStudy (marmar10):

help! prove cos^2 x/1-sinx=1+sinx

OpenStudy (callisto):

\[\frac{\cos^2x}{1-\sin x}\]Multiply the expression by the conjugate \(\frac{1+\sin x}{1+\sin x}\), then we get \[\frac{\cos^2x (1+\sin x)}{(1-\sin x)(1+\sin x)}\]Expand the denominator, what do you get?

OpenStudy (marmar10):

thank you!!!

OpenStudy (callisto):

You know how to proceed?

OpenStudy (marmar10):

how do you get rid of the cosx?

OpenStudy (callisto):

What do you get when you expand the expression in the denominator?

OpenStudy (callisto):

Hint: Use \((a+b)(a-b) = a^2-b^2\) to expand \((1+\sin x)(1-\sin x)\)

OpenStudy (marmar10):

ohh wow haha I got it now. thanks